OpenStudy (anonymous):

Suppose that g is a quadratic function such that g(x) >= 0 only on the interval (-1,7) and g(1)=36. a. Find the rule of g(x) b. Find all Valuse of x such that g(x)=10. c. Find the maximum value of g(x) or explain why no such value exists.

OpenStudy (anonymous):

Since \(g\) is quadratic, its general form is\[g(x)=ax^2+bx+c\] Since \(g\ge0\) only on the provided interval, you're essentially told that \(g=0\) when \(x=-1\) and \(x=7\). This means you can factor the above form to look like \[g(x)=a(x+1)(x-7)\] Given that \(g(1)=36\), you have \[a(1+1)(1-7)=36~~\iff~~a=-3\]